# Balanced and Elastic End-to-end Training of Dynamic LLMs

### v1

@@ Abstract.
We propose an autonomous dynamic load balancing solution, DynMo, which provably achieves maximum reduction in workload imbalance and adaptively equalizes compute loads across workers in pipeline-parallel training. In addition, DynMo dynamically consolidates computation onto fewer workers without sacrificing training throughput, allowing idle workers to be released back to the job manager. DynMo supports both single-node multi-GPU systems and multi-node GPU clusters, and can be used in practical deployment. Compared to static distributed training solutions such as Megatron-LM and DeepSpeed, DynMo accelerates the end-to-end training of dynamic GPT models by up to 1.23× for MoEs, 3.18× for parameter pruning, 2.23× for layer freezing, 4.02× for sparse attention, 4.52× for early exit, and 1.17× for MoDs. DynMo is available at https://anonymous.4open.science/r/DynMo-4D04/.
conference: The International Conference for High Performance Computing, Networking, Storage, and Analysis; November 16–21; Somewhereisbn: doi: copyright: none
@@ 1. Introduction
Figure 1. Average idleness percentage of GPUs (per iteration) for training dynamic GPT models (Radford et al., 2018) on 720 H100 GPUs in total, excluding MoEs which uses 128 H100 GPUs in total. All use models parameterized to have between 24 and 48 layers, and except for MoEs, which we report the average idleness percentage for Mixtral 8x7b and LlaMA-MoE-3.5B models. For pipeline parallelism, we use the highest performing (single) pipeline parallelism scheme known to the authors: the ”almost zero-bubble pipeline parallelism” scheme (Qi et al., 2024). The bubble ratios are measured on a hybrid of pipeline and data parallelism. 1 MoE: we observe $\sim$25bubble ratio in the pipeline on Mixtral 8x7b (Jiang et al., 2024) and LLaMA-MoE-3.5b (Team, 2024), arising from the load imbalance imposed by the routing schemes used in token choice (S-BASE (Lewis et al., 2021a) and load imbalance with auxiliary loss (Jiang et al., 2024)). 2 Gradual pruning of model parameters: we observe almost a five fold increase in idleness at 90% sparsity levels. Note that idleness at Dense is the inherent pipeline bubbles of a static model. 3 Layer freezing: SoTA freezing schemes that incorporate load balancing (Egeria (Wang et al., 2022) and AutoFreeze (Liu et al., 2021)) yield $\sim$40bubble ratio. 4 Dynamic Sparse Flash Attention: locality sensitive hashing with support for flash attention (Pagliardini et al., 2023) exhibits a 4x increase in the bubble ratio over the baseline dense attention. 5 Early exit: SoTA early exit methods (CALM (Schuster et al., 2022) and ADP-C (Liu et al., 2022b)) exhibits up to 5x increase in the bubble ratio over the baseline (w/o early exit), mainly due to the accumulation of bubbles in late layers. 6 MoD: we observe $\sim$18bubble ratio in the pipeline, arising from the load imbalance imposed by the routing scheme of expert choices that lacks information about future tokens (Raposo et al., 2024).
@@
In traditional LLM training, each pipeline stage has a fixed workload known in advance. In contrast, emerging schemes designed to reduce computational demands introduce dynamic workloads. These include: neural networks where different input samples take different pathways through the model layers, e.g. gated neural networks (Shazeer et al., 2017a), sparsely activated Mixture of Experts (MoEs) (Zhou et al., 2022b), Switch Transformers (Fedus et al., 2022), Mixture of Depths (MoDs) (Raposo et al., 2024) etc., gradual pruning where the parameters of a model are pruned (i.e. sparsified) during training (Gale et al., 2019), freeze training where some of the layers of the model are adapatively frozen during training (Wang et al., 2022), different schemes to dynamically sparsify the attention matrix (Liu et al., 2022a; Pagliardini et al., 2023; Tay et al., 2020), and early exit strategies where tokens skip remaining layers based on an exit decision (Elbayad et al., 2020; Schuster et al., 2022; Liu et al., 2022b; Kim et al., 2022). Beyond computational efficiency, dynamic models are also explored for benefits like improved generalization and explainability. We refer readers to (Han et al., 2021; Tay et al., 2022) for surveys on dynamic modeling.
@@
One of the main downsides of using dynamic models is the load imbalance they introduce in pipeline parallelism, which significantly reduces LLM training throughput (Zhou et al., 2022a; He et al., 2022). For example, Figure 1 shows the average GPU idleness for GPT models with varying depths and different dynamic model types. This imbalance appears as bubbles in the pipeline, where an accelerator stalls while waiting for delayed work from its neighboring stage. As illustrated in the figure, idleness can range from 18% to 5x, depending on the scheme e.g., layer freezing introduces up to 40% idleness for a 40-layer model. Note that these bubbles are distinct from the inherent gaps in pipeline scheduling, such as idle periods between forward and backward passes in GPipe (Huang et al., 2019). The former, which stems from dynamic workloads, is the primary focus of this paper.
@@
Innovative designs of dynamic models aim to reduce computational cost, but without effective load balancing, their benefits practically fail to translate into actual performance gains during distributed training (Cheng et al., 2024). To address this gap, we introduce DynMo, an elastic load-balancing framework tailored for dynamic models. It is the first work to study pipeline stalls caused by training dynamic models. DynMo ensures balanced pipeline stages by dynamically redistributing workloads across accelerators whenever imbalance arises, thereby improving computational efficiency and reducing training costs. The framework incorporates two different dynamic balancers, both proven to converge to the optimal workload balance among workers. Our experiments demonstrate that DynMo incurs negligible overhead and scales effectively in both single-node multi-GPU setups and multi-node environments commonly used for large-scale LLM training with hybrid parallelism.
@@
  • We introduce DynMo, which supports dynamic models and significantly improves their end-to-end training efficiency, making them practical for real-world use.
  • We invoke DynMo to rebalance at regular intervals without prior knowledge of dynamism, hence balancing the load in a fully automated and transparent manner. DynMo is orthogonal to the underlying pipeline and dynamism schemes, ensuring compatibility with diverse dynamic compute and model reduction methods.
@@ 2. Dynamic Models
In the remainder of this section, we build on the formal load balancing definition by extending the model to capture the dynamism introduced by each of our six example cases. An elaborate formulation of the dynamism arising from each of the six schemes will be provided as supplementary material.In each of the six cases, we define the load on each worker $N_{j}$ to calculate the load imbalance $\Delta L^{(k)}$ at time step $k$. Ideally, $L_{j}^{(k)}\approx L_{j^{\prime}}^{(k)}$ for all workers, but the dynamic nature of the six cases leads to an imbalance. Defining maximum and minimum loads:
@@ 2.1. Mixture of Experts
In MoEs (Shazeer et al., 2017b), input tokens are routed to specialized sub-networks (experts) instead of a single feed-forward network, improving efficiency but introducing load imbalance when those experts are distributed among workers. Empirical results from Mixtral 8x7B (Jiang et al., 2024) show up to 25% imbalance. Prior approaches attempt mitigation via auxiliary losses (Jiang et al., 2024; Lewis et al., 2021a), but routing remains suboptimal. Recent strategies, such as DeepSeek V3 (DeepSeek-AI et al., 2025), introduce bias terms in token-to-expert affinity scores (Wang et al., 2024), achieving up to  8% imbalance per layer in a 3B model. However, imbalance compounds across layers and is expected to be higher in larger models. Load imbalance propagates through training, creating pipeline stalls, especially during all_to_all communication. Addressing this remains crucial for efficient MoE training in distributed systems.
Let =$\mathcal{E}=\{e_{1},e_{2},\dots,e_{k}\}$ be the set of experts, each assigned to a worker $\mathcal{N}_{j}\in\mathcal{N}$. A routing function $R:\mathcal{T}\to\mathcal{N}$ maps tokens to workers. The load of a worker $L_{j}$ aggregating the workload $c_{e_{j}}$ all tokens it is assigned, at time step $k$:
@@ 2.2. Parameter Pruning
Let =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers, each assigned to a worker $\mathcal{N}_{j}$. During distributed training, workload varies as parameters are pruned. The fraction of retained parameters in layer $l_{i}$ at time step $k$ is $p_{i}^{(k)}$, the total network compute being $C^{(k)}$, where $A^{(k)}:\mathcal{L}\to\mathcal{N}$ be the layer-to-worker assignment. The load on worker $\mathcal{N}_{j}$ is:
@@
The load per worker defined above can be used with Equations 1 and 2 to get the overall system imbalance introduced by the pipeline stalls occurring while some workers wait for heavily pruned layers to complete (Zhu and Gupta, 2017; Frankle and Carbin, 2018; Bellec et al., 2017). Supplementary material discusses pruning criteria, structure, and scheduling.
@@ 2.3. Layer Freezing
Let =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers in a DNN, each assigned to a worker $\mathcal{N}_{i}$. Earlier layers often converge faster (Wang et al., 2022), enabling their freezing to save computation. Let $f_{i}^{(k)}\in\{0,1\}$ indicate whether layer $l_{i}$ is frozen at time step $k$:
@@ 2.4. Dynamic Sparse Flash Attention
Let =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers, each using dynamic sparse attention. The workload depends on the sparsity level, which changes dynamically. Let $s_{i}^{(k)}\in[0,1]$ be the sparsity factor of layer $l_{i}$ at time $k$, representing the fraction of nonzero elements in the attention matrix. For $c_{i}$ being the initial workload of layer $l_{i}$. Let $A^{(k)}:\mathcal{L}\to\mathcal{N}$ be the layer-to-worker assignment. The load on worker $\mathcal{N}_{j}$ is: at time $k$ is:
@@ 2.5. Early Exit
Let =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers where tokens can exit early. Let $t^{(k)}$ be the total tokens at time $k$, and $t_{i}^{(k)}$ as the tokens processed by layer $l_{i}$. Typically, $t_{i}^{(k)}$ decreases with depth (Schuster et al., 2022; Liu et al., 2022b). For $c_{i}$ being the initial workload. For $c_{i}$ being the initial workload, the total workload of all layers is:
@@ 2.6. Mixture of Depths
Let =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers, with $t^{(k)}$ as the total number of tokens at time $k$, and $t_{i}^{(k)}$ as the tokens processed by layer $l_{i}$. MoD enables tokens to skip layers based on expert predictions, leading to fluctuating workloads. The routing weight $r_{i}^{(k)}$ determines whether tokens bypass a layer, and the total effective compute of $l_{i}$ at time $k$ is:
@@ 3.1. Overview
DynMo takes as input a model, the number of training iterations, and several dynamism configuring arguments (that have default values if the user wishes to use DynMo out-of-the-box). We start the training with the original model and train it until the dynamism to apply is reached. The frequency varies by the target case, in layer freezing it is as frequent as every 50 iterations, while in parameter tuning the pruning frequency is in the range of 3000-7000 iterations. The model or control flow is altered only if the training is in the dynamism stage. In this stage, the model is adjusted by either being modified or the control flow inside the model changes, until it meets the training stopping criteria. The first iteration after each dynamism operation is used for profiling the time it takes to execute each layer in the altered model and the memory usage of all workers (accelerators) in the pipeline. Next, DynMo collects the profiling information and decides on balancing the workload by moving layers across pipeline stages based on the execution times of individual layers to minimize the pipeline stalls, subject to the constraints of memory capacity per worker. DynMo also attempts to re-pack the total workload into fewer number of GPUs if the re-packing feature is enabled by the end user. Once the training is out of the dynamism region, the balanced pipeline continues to execute with the model.
@@ 3.3. Load Balancing
We demonstrate that the two load balancing schemes meet the goals for optimal load balancing by using the following lemmas. Due to the space limitation, formal proofs are omitted, and will be included in the supplementary material.

### v2

@@ Abstract.
We propose an autonomous dynamic load balancing solution, DynMo, which provably achieves maximum reduction in workload imbalance and adaptively equalizes compute loads across workers in pipeline-parallel training. In addition, DynMo dynamically consolidates computation onto fewer workers without sacrificing training throughput, allowing idle workers to be released back to the job manager. DynMo supports both single-node multi-GPU systems and multi-node GPU clusters, and can be used in practical deployment. Compared to static distributed training solutions such as Megatron-LM and DeepSpeed, DynMo accelerates the end-to-end training of dynamic GPT models by up to 1.23x for MoEs, 3.18x for parameter pruning, 2.23x for layer freezing, 4.02x for sparse attention, 4.52x for early exit, and 1.17x for MoDs.
Large Language Models, Load Balancing, Pipeline Parallelism
conference: The International Conference for High Performance Computing, Networking, Storage, and Analysis; November 16–21; St. louis, MOccs: Computer systems organization Parallel architectures
@@ 1. Introduction
Figure 1. Average idleness percentage of GPUs (per iteration) for training dynamic GPT models (Radford et al., 2018) on 720 H100 GPUs in total, excluding MoEs which uses 128 H100 GPUs in total. We use models parameterized to have between 24 and 48 layers, except for the MoEs user case for which we report the average idleness percentage for Mixtral 8x7b and LlaMA-MoE-3.5B models. For pipeline parallelism, we use the highest performing (single) pipeline parallelism scheme known to the authors: the ”almost zero-bubble pipeline parallelism” scheme (Qi et al., 2024). The bubble ratios are measured on a hybrid of pipeline and data parallelism. 1 MoE: we observe $\sim$25% bubble ratio in the pipeline on Mixtral 8x7b (Jiang et al., 2024) and LLaMA-MoE-3.5b (Team, 2024), arising from the load imbalance imposed by the routing schemes used in token choice (S-BASE (Lewis et al., 2021a) and load imbalance with auxiliary loss (Jiang et al., 2024)). 2 Gradual pruning of model parameters: we observe almost a five fold increase in idleness at 90% sparsity levels. Note that idleness at Dense is the inherent pipeline bubbles of a static model. 3 Layer freezing: SoTA freezing schemes that incorporate load balancing (Egeria (Wang et al., 2022) and AutoFreeze (Liu et al., 2021)) yield $\sim$40% bubble ratio. 4 Dynamic Sparse Flash Attention: locality sensitive hashing with support for flash attention (Pagliardini et al., 2023) exhibits a 4x increase in the bubble ratio over the baseline dense attention. 5 Early exit: SoTA early exit methods (CALM (Schuster et al., 2022) and ADP-C (Liu et al., 2022b)) exhibits up to 5x increase in the bubble ratio over the baseline (w/o early exit), mainly due to the accumulation of bubbles in late layers. 6 MoD: we observe $\sim$18% bubble ratio in the pipeline, arising from the load imbalance imposed by the routing scheme of expert choices that lacks information about future tokens (Raposo et al., 2024).
@@
In traditional LLM training, each pipeline stage has a fixed workload known in advance. In contrast, emerging schemes designed to reduce computational demands introduce dynamic workloads. These include: neural networks where different input samples take different pathways through the model layers, e.g. gated neural networks (Shazeer et al., 2017a), sparsely activated Mixture of Experts (MoEs) (Zhou et al., 2022b), Switch Transformers (Fedus et al., 2022), Mixture of Depths (MoDs) (Raposo et al., 2024) etc., gradual pruning where the parameters of a model are pruned (i.e. sparsified) during training (Gale et al., 2019), freeze training where some of the layers of the model are adapatively frozen during training (Wang et al., 2022), different schemes to dynamically sparsify the attention matrix (Liu et al., 2022a; Pagliardini et al., 2023; Tay et al., 2020), and early exit strategies where tokens skip remaining layers based on an exit decision (Elbayad et al., 2020; Schuster et al., 2022; Liu et al., 2022b; Kim et al., 2022). Beyond computational efficiency, dynamic models are also explored for benefits like improved generalization and explainability. We refer readers to (Han et al., 2021; Tay et al., 2022) for surveys on dynamic LLMs.
@@
One of the main downsides of using dynamic models is the load imbalance they introduce in pipeline parallelism, which significantly reduces LLM training throughput (Zhou et al., 2022a; He et al., 2022). For example, Figure 1 shows the average GPU idleness for GPT models with varying depths and different dynamic model types. This imbalance appears as bubbles in the pipeline, where an accelerator stalls while waiting for work to be passed from its neighboring worker. As illustrated in the figure, idleness can range from 18% to 5x, depending on the scheme, e.g., layer freezing introduces up to 40% idleness for a 40-layer model. Note that these bubbles are distinct from the inherent gaps in pipeline scheduling, such as idle periods in the pipeline wind-up and wind-down. The former, which stems from dynamic workloads, is the primary focus of this paper.
@@
Innovative designs of dynamic models aim to reduce computational cost, but without effective load balancing, their benefits practically fail to translate into actual performance gains during distributed training (Cheng et al., 2024). To address this gap, we introduce DynMo, an elastic load-balancing framework tailored for dynamic models. It is the first work to study pipeline stalls caused by training dynamic models. DynMo ensures balanced pipeline stages by dynamically redistributing workloads across accelerators whenever imbalance arises, thereby improving computational efficiency and reducing training costs. The framework incorporates two different dynamic balancers, both proven to converge to the optimal workload balance among workers. Our experiments demonstrate that DynMo incurs negligible overhead and scales effectively in both single-node multi-GPU and multi-node multi-GPU setups.
@@
  • We introduce DynMo, a framework for load balancing dynamic LLMs to improves their end-to-end training efficiency, making them practical for real-world use. We invoke DynMo to rebalance at regular intervals without prior knowledge of dynamism, hence balancing the load in a fully automated and transparent manner. DynMo is orthogonal to the underlying pipeline and dynamism schemes, ensuring compatibility with diverse dynamic compute and model reduction methods.
@@ 2. Dynamic Models
In the remainder of this section, we build on the formal load balancing definition by extending the model to capture the dynamism introduced by each of our six example cases. In each of the six cases, we define the load on each worker $N_{j}$ to calculate the load imbalance $\Delta L^{(k)}$ at time step $k$. Ideally, $L_{j}^{(k)}\approx L_{j^{\prime}}^{(k)}$ for all workers, but the dynamic nature of the six cases leads to an imbalance. Defining maximum and minimum loads:
@@ 2.1. Mixture of Experts
In MoEs (Shazeer et al., 2017b), input tokens are routed to specialized sub-networks (experts) instead of a single feed-forward network, improving efficiency but introducing load imbalance when those experts are distributed among workers. Empirical results from Mixtral 8x7B (Jiang et al., 2024) show up to 25% imbalance. Prior approaches attempt mitigation via auxiliary losses (Jiang et al., 2024; Lewis et al., 2021a), yet routing remains suboptimal. Recent strategies, such as DeepSeek V3 (DeepSeek-AI et al., 2025), introduce bias terms in token-to-expert affinity scores (Wang et al., 2024), achieving up to  8% imbalance per layer in a 3B model. However, imbalance compounds across layers and is expected to be higher in larger models. Load imbalance propagates through training, creating pipeline stalls, especially during all_to_all communication. Addressing this remains crucial for efficient MoE training in distributed systems.
Let $\mathcal{E}=\{e_{1},e_{2},\dots,e_{k}\}$ be the set of experts, each assigned to a worker $\mathcal{N}_{j}\in\mathcal{N}$. A routing function $R:\mathcal{T}\to\mathcal{N}$ maps tokens to workers. The load of a worker $L_{j}$ aggregating the workload $c_{e_{j}}$ all tokens it is assigned, at time step $k$:
@@ 2.2. Parameter Pruning
Let $\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers, each assigned to a worker $\mathcal{N}_{j}$. During distributed training, workload varies as parameters are pruned. The fraction of retained parameters in layer $l_{i}$ at time step $k$ is $p_{i}^{(k)}$, the total network compute being $C^{(k)}$, where $A^{(k)}:\mathcal{L}\to\mathcal{N}$ be the layer-to-worker assignment. The load on worker $\mathcal{N}_{j}$ is:
@@
The load per worker defined above can be used with Equations 1 and 2 to get the overall system imbalance introduced by the pipeline stalls occurring while some workers wait for heavily pruned layers to complete (Zhu and Gupta, 2017; Frankle and Carbin, 2018; Bellec et al., 2017).
@@ 2.3. Layer Freezing
Let $\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers in an LLM, each assigned to a worker $\mathcal{N}_{i}$. Earlier layers often converge faster (Wang et al., 2022), enabling their freezing to save computation. Let $f_{i}^{(k)}\in\{0,1\}$ indicate whether layer $l_{i}$ is frozen at time step $k$:
@@ 2.4. Dynamic Sparse Flash Attention
Let $\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers, each using dynamic sparse attention. The workload depends on the sparsity level, which changes dynamically. Let $s_{i}^{(k)}\in[0,1]$ be the sparsity factor of layer $l_{i}$ at time $k$, representing the fraction of nonzero elements in the attention matrix. For $c_{i}$ being the initial workload of layer $l_{i}$. Let $A^{(k)}:\mathcal{L}\to\mathcal{N}$ be the layer-to-worker assignment. The load on worker $\mathcal{N}_{j}$ is: at time $k$ is:
@@ 2.5. Early Exit
Let $\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers where tokens can exit early. Let $t^{(k)}$ be the total tokens at time $k$, and $t_{i}^{(k)}$ as the tokens processed by layer $l_{i}$. Typically, $t_{i}^{(k)}$ decreases with depth (Schuster et al., 2022; Liu et al., 2022b). For $c_{i}$ being the initial workload. For $c_{i}$ being the initial workload, the total workload of all layers is:
@@ 2.6. Mixture of Depths
Let $\mathcal{L}=\{l_{1},l_{2},\dots,l_{d}\}$ be the set of layers, with $t^{(k)}$ as the total number of tokens at time $k$, and $t_{i}^{(k)}$ as the tokens processed by layer $l_{i}$. MoD enables tokens to skip layers based on expert predictions, leading to fluctuating workloads. The routing weight $r_{i}^{(k)}$ determines whether tokens bypass a layer, and the total effective compute of $l_{i}$ at time $k$ is:
@@ 3.1. Overview
DynMo takes as input a model, the number of training iterations, and several dynamism configuring arguments (that have default values if the user wishes to use DynMo out-of-the-box). We start the training with the original model and train it until the dynamism to apply is reached. The frequency varies by the target case, in layer freezing it is as frequent as every 50 iterations, while in parameter tuning the pruning frequency is in the range of 3000-7000 iterations. The model or control flow is altered only if the training is in the dynamism stage. In this stage, the model is adjusted by either being modified or when the control flow inside the model changes, until it meets the training stopping criteria. The first iteration after each dynamism operation is used for profiling the time it takes to execute each layer in the altered model and the memory usage of all workers (accelerators) in the pipeline. Next, DynMo collects the profiling information and decides on balancing the workload by moving layers across pipeline stages based on the execution times of individual layers to minimize the pipeline stalls, subject to the constraints of memory capacity per worker. DynMo also attempts to re-pack the total workload into fewer number of GPUs if the re-packing feature is enabled by the end user. Once the training is out of the dynamism region, the balanced pipeline continues to execute with the model.
@@ 3.3. Load Balancing
We demonstrate that the two load balancing schemes meet the goals for optimal load balancing by using the following lemmas.
@@
Proof. We will prove by contradiction. Suppose a centralized load balancer $L_{c}$ over $\mathcal{N}$ workers satisfies maximum reduction in the imbalance $\mathcal{N}_{i}$ when $\mathcal{N}_{i}$ has a bubble ratio higher then the minimum. By the definition of maximum reduction in load balance, $L_{c}$ must preserve minimum differential between the loads of workers $\mathcal{N}_{i}$ and $\mathcal{N}_{j}$, which $\mathcal{N}_{i}$ and $\mathcal{N}_{j}$ have the minimum load and maximum loads in $\mathcal{N}$, respectively. Consequently, increasing the bubble ratio of $\mathcal{N}_{j}$ changes the load difference between $\mathcal{N}_{i}$ and $\mathcal{N}_{j}$. This is in contradictory of $L_{c}$ achieving maximum reduction on imbalance. ∎
@@
Proof. We leverages core ideas from Lyapunov optimization. We first define a potential function, $\phi$, that measures at each round the total magnitude of workload gaps in the system:
$\forall r\geq 0:\phi(r)=\sum_{u,v\in V}\left|x_{u}(r)-x_{v}(r)\right|$
Similar to a Lyapunov function, $\phi$ maps the system state (in this case, a vector of workloads for $\mathcal{N}$ workers) at any given round to a non-negative scalar value that describes the desirability of the current system state. As $\phi$ decreases toward 0, the system state becomes more desirable; i.e. the workload is balanced across $\mathcal{N}$. As in a standard Lyapunov optimization, we show below that the modifications to a system state caused by executing a single round of our max neighbor algorithm will drift the value of $\phi$ toward zero in a non-decreasing manner. We establish a probabilistic lower bound for the amount of drift in a given round to obtain our time bounds.
[…]
For a maximum number of rounds to converge to the minimum imbalance:
$O\left(\mathrm{min}\left\{\mathcal{N}^{2}\mathrm{log}\left(\frac{S\mathcal{N}}{\gamma}\right)\mathrm{log}~\mathcal{N},\frac{SN~\mathrm{log}~\mathcal{N}}{\gamma}\right\}\right)$
Note that these two bounds essentially coincide at $\tilde{O}(\mathcal{N}^{2})$ with $\gamma=\Theta(S/n)$, where the notation $\tilde{O}$ hides logarithmic factors. In other words, if we want all nodes to have the same workload up to a constant factor, the max neighbor strategy uses $\tilde{O}(\mathcal{N}^{2})$ rounds. We first note that if we arrive at a round r in which $\phi(r)\leq\gamma$, then the system ends this round $\gamma$-converged, i.e. the sum of the gaps is at most $\gamma$, and thus clearly any individual gap is at most $\gamma$. Since $\phi$ is monotonically non-increasing, it follows that every round ${r}^{\prime}\geq r$ is also $\gamma$-converged. So we just need to show that with high probability, $\phi$ will decrease to $\gamma$ in the time bound stated by the theorem statement.
For each $r\geq 1$, we call r “good” if and only if $\phi(r-1)-\phi(r)\geq s_{max}(r-1)/(60~ln(2n))$. We next calculate how many good rounds guarantee that $\phi$ falls below $\gamma$. To do so, we first note that, non-good rounds cannot increase $\phi$, so we are safe to focus only on reductions generated by good rounds in calculating our bound.
By the definition of $\phi$, for each $r\geq 1$ we know that $\phi(r)<s_{max}(r)n^{2}$. It follows that if $r$ is a good round, then it decreases $\phi(r-1)$ by a multiplicative factor less than $(1-\frac{1}{60n^{2}ln(2n)})$. Finally, we also observe that $s_{max}(0)\leq S$ and therefore $\phi(0)<Sn^{2}$. Leveraging these observations, to find the number of good rounds needed to decrease $\phi$ below $\gamma$, we just need to find the minimum $s$ time steps such that
$Sn^{2}\left(1-\frac{1}{60n^{2}ln(2n)}\right)\leq\gamma$
A simple calculation implies that $s_{con}=60n^{2}ln(2n)ln(Sn^{2}\gamma^{-1})$ is sufficient to satisfy this inequality. We have now established that after $s_{con}$ good rounds the system will be $\gamma$-converged for all future rounds. We are left to bound the number of rounds required to generate $s_{con}$ good rounds with high probability.
For each round $r$, let $X_{r}$ be the random indicator variable that evaluates to $1$ if round $r$ is good and otherwise evaluates to 0. We know a given round $r$ is good with probability at least $1/\mathcal{N}$, regardless of the history of the execution through the round r1r-1. We cannot, however, directly leverage this observation to calculate (and concentrate) the expected sum of $X$ variables for a given execution length, as the distribution determining a given $X_{r}$ might depend in part on the outcome of previous experiments. To overcome this issue, we define for each round $r$, a trivial random indicator variable $\hat{X}_{r}$ that evaluates to $1$ with independent probability $1/N$ and otherwise evaluates to 0. Note that for each $r$, $X_{r}$ stochastically dominates $\hat{X}_{r}$, and therefore for each $s>0,Y_{s}=\sum_{r=1}^{s}X_{r}$ stochastically dominates $s>0,\hat{Y}_{s}=\sum_{r=1}^{s}\hat{X}_{r}$. It follows for any $s>0$, if $\hat{Y}_{s}\geq s_{con}$ with some probability $p$ then $Y_{s}\geq s_{con}$ with probability at least $p$.
A Chernoff bound applied to $\hat{Y}_{s}$, for $s=c.s_{con}$ (where $c\geq 1$ is a sufficiently large constant defined with respect to the constants in $s_{con}$ and the constants in the Chernoff form used), provides that $\hat{Y}_{s}\geq s_{con}$ with high probability, and therefore so is $Y_{s}$. To conclude the proof, we note that $c.s_{con}\in O\left(\mathcal{N}^{2}log(\frac{S\mathcal{N}}{\gamma})log\mathcal{N}\right)$, as required by the theorem $\gamma$ statement. ∎